# OrbitZoo: Multi-Agent Reinforcement Learning Environment for Orbital Dynamics

### v1

@@ Appendix 0.G Data and residual analysis
Figures 22 through 25 compare the predictions of a three-day propagation starting on November 27th at around 5:00 AM and ending three days later. The aggregation of 10 consecutive ephemeris files allows for the residual analysis of the initial three-day prediction (from the first ephemeris file) and the subsequent predictions made in the following 9 ephemeris files.
Figures  26 through 29 show the evolution of the three positional vectors for four different satellites. Since the plotted position is derived from a combination of ephemeris files, using only the most accurate segments from each file, a discontinuity occurs at the transitions between files when one ends and another begins. As expected, the position is oscillatory in nature, reaching its maximum value when the satellite is at apogee (the farthest point from Earth along the major axis of an elliptical orbit) and its minimum value when the satellite is at perigee (the closest point to Earth along the same axis).
@@
A propagation of 16.6 hours (1000 steps) was performed using Orekit, Figures 34 through 37 display the residuals of this propagation for four different satellites.
@@
First, limited information on how Starlink produces its ephemeris data can lead to inconsistencies when attempting to approximate the satellites’ trajectories. Researchers, [32] suggest that Starlink may use multiple propagators and even switch among them mid-propagation, introducing further uncertainty into a single ephemeris file. Consequently, reconstructing each satellite’s real orbital trajectory from Starlink’s ephemeris data may vary across different satellites.

### v2

@@ Appendix H Data and residual analysis
Figures 30 through 33 compare the predictions of a three-day propagation starting on November 27th at around 5:00 AM and ending three days later. The aggregation of 10 consecutive ephemeris files allows for the residual analysis of the initial three-day prediction (from the first ephemeris file) and the subsequent predictions made in the following 9 ephemeris files.
Figures  34 through 37 show the evolution of the three positional vectors for four different satellites. Since the plotted position is derived from a combination of ephemeris files, using only the most accurate segments from each file, a discontinuity occurs at the transitions between files when one ends and another begins. As expected, the position is oscillatory in nature, reaching its maximum value when the satellite is at apogee (the farthest point from Earth along the major axis of an elliptical orbit) and its minimum value when the satellite is at perigee (the closest point to Earth along the same axis).
@@
A propagation of 16.6 hours (1000 steps) was performed using Orekit, Figures 42 through 45 display the residuals of this propagation for four different satellites.
@@
First, limited information on how Starlink produces its ephemeris data can lead to inconsistencies when attempting to approximate the satellites’ trajectories. Researchers [44] suggest that Starlink may use multiple propagators and even switch among them mid-propagation, introducing further uncertainty into a single ephemeris file. Consequently, reconstructing each satellite’s real orbital trajectory from Starlink’s ephemeris data may vary across different satellites.
@@
Appendix I Broader Impacts
This work advances RL research in high-dimensional, discrete and continuous control domains relevant to space operations, incorporating perturbative forces, model uncertainties, and ephemeris-based validation. Notably, OrbitZoo includes support for federated learning, enabling decentralized training across multiple agents or environments while preserving data locality, an important step toward scalable and privacy-aware coordination in space missions.
Potential positive impacts include increased autonomy and efficiency in satellite operations, reduced reliance on ground-based control, and more resilient responses to unexpected events like debris conjunctions. OrbitZoo can accelerate the development of safe, adaptive control strategies for real-world missions, and democratize access to space-focused RL research by offering an open, extensible platform.
Potential negative impacts include over-reliance on black-box models that may behave unpredictably in safety-critical situations. The deployment of learned policies without robust interpretability or certification frameworks could lead to unintended behaviors or mission failures. Additionally, advances in autonomous maneuvering might be misused in military or competitive commercial contexts without adequate regulatory oversight.
Our experiments demonstrate that policies trained in OrbitZoo can closely replicate real-world satellite behavior, including Starlink trajectories, positioning the platform as a benchmark for trustworthy and autonomous space systems.